# MonkeyOCR v1.5 Technical Report: Unlocking Robust Document Parsing for Complex Patterns

### v1

@@ Abstract
Document parsing is a core task in document intelligence, supporting applications such as information extraction, retrieval-augmented generation, and automated document analysis. However, real-world documents often feature complex layouts with multi-level tables, embedded images or formulas, and cross-page structures, which remain challenging for existing OCR systems. We introduce MonkeyOCR v1.5, a unified vision–language framework that enhances both layout understanding and content recognition through a two-stage parsing pipeline. The first stage employs a large multimodal model to jointly predict document layout and reading order, leveraging visual information to ensure structural and sequential consistency. The second stage performs localized recognition of text, formulas, and tables within detected regions, maintaining high visual fidelity while reducing error propagation. To address complex table structures, we propose a visual consistency–based reinforcement learning scheme that evaluates recognition quality via render-and-compare alignment, improving structural accuracy without manual annotations. Additionally, two specialized modules, Image-Decoupled Table Parsing and Type-Guided Table Merging, are introduced to enable reliable parsing of tables containing embedded images and reconstruction of tables crossing pages or columns. Comprehensive experiments on OmniDocBench v1.5 demonstrate that MonkeyOCR v1.5 achieves state-of-the-art performance, outperforming PPOCR-VL and MinerU 2.5 while showing exceptional robustness in visually complex document scenarios.
@@ 1 Introduction
Traditional pipeline-based approaches [28; 6] break down document parsing into a series of subtasks, such as layout detection, text and formula detection, text recognition, and table or formula recognition, with each handled by a dedicated model. This multi-stage process is prone to error accumulation, where mistakes in earlier steps propagate and compromise overall performance. In contrast, end-to-end models [1; 3; 19; 8] process the entire document image in a single pass; however, the high resolution of document images produces a massive number of visual tokens, and the quadratic complexity of self-attention mechanisms creates a significant computational bottleneck. To address these limitations, MonkeyOCR proposed the SRR paradigm, which decouples document parsing into structure detection, content recognition, and reading order prediction. This design streamlines the conventional multi-stage pipeline, effectively mitigating cumulative errors while avoiding the substantial computational overhead associated with full-page end-to-end processing, thereby advancing intelligent multimodal document understanding (Fig. 2). Mineru 2.5 [18] further simplify the three-stage framework by employing a unified large multimodal model to jointly predict document layout and reading order, followed by content recognition. PPOCR-VL [5] adopts a similar three-stage methodology, utilizing lightweight models for structural analysis and reading order prediction, and subsequently applying a large multimodal model for content recognition.
@@
Experimental results show that MonkeyOCR v1.5 achieves state-of-the-art performance on the widely adopted OmniDocBench v1.5 benchmark, surpassing the previous best-performing methods PPOCR-VL and MinerU 2.5 by 1.0% and 2.2%, respectively, in overall performance. Notably, MonkeyOCR v1.5 demonstrates stronger robustness in complex scenarios. On newspaper documents characterized by dense text and intricate layouts, it achieves the best performance among all competitors. Moreover, on the OCRFlux-complex dataset, our method significantly outperforms the previous state-of-the-art PPOCR-VL by 8.2%. Beyond accuracy improvement, MonkeyOCR v1.5 further extends its capability to handle embedded image recovery, cross-page table reconstruction, and multi-column table merging, showcasing strong generalization in complex real-world document parsing scenarios.
@@
We propose MonkeyOCR v1.5, a vision-language-based document parsing framework designed for robust and efficient OCR in complex, real-world layouts. Compared with the previous version, v1.5 simplifies the pipeline into two stages and introduces a visual-consistency-based reinforcement learning paradigm that improves table recognition accuracy without relying on large-scale manual annotations. In addition, to address challenges such as embedded images and cross-page or cross-column table merging that other methods struggle with (Tab. 1), we incorporate an image-decoupled and type-guided table recognition module. Together, these advancements make MonkeyOCR v1.5 a scalable and high-fidelity OCR system well-suited for heterogeneous document understanding.
@@
MinerU2.5[18] × × × × ×
MonkeyOCR[13] × × × × × × ×
OCRFlux[2] × × × × ×
dotsOCR[25] × × × × × × ×
PaddleOCR-VL[5] × × × × × ×
@@ 2.1 Overall Pipeline
The overall pipeline of MonkeyOCR v1.5 consists of two sequential yet lightweight stages: layout analysis and content recognition, as described in Fig. 3. This design balances global structural understanding and fine-grained semantic parsing while maintaining computational efficiency.
@@ 2.3 Image-Decoupled Table Parsing
As illustrated in Fig. 5, we first employ YOLOv10 [26] to detect embedded images within table regions. Each detected figure is replaced by a precisely sized placeholder mask, while a one-to-one mapping between placeholder IDs and cropped image files is maintained. The masked table is then passed to the recognizer, which is trained with an auxiliary objective encouraging it to treat placeholders as atomic tokens and to produce an intermediate HTML representation containing <img> tags at the corresponding locations. During post-processing, each <img> tag is deterministically replaced with its associated image according to the stored mapping, yielding a visually complete table while preserving clean textual outputs. This decoupled design effectively separates image restoration from structural recognition, improving both text accuracy and visual fidelity in complex tables.
@@ 2.4 Type-Guided Table Merging
PaddleOCR-VL [5] 91.9 0.039 88.7 91.0 94.9 0.048
MinerU2.5 [18] 90.7 0.047 88.5 88.2 92.4 0.044
MonkeyOCR-pro-3B [13] 88.9 0.075 87.3 86.8 90.6 0.128
dots.ocr [25] 88.4 0.048 83.2 86.8 90.6 0.053
MonkeyOCR-3B 87.1 0.075 87.5 81.4 85.9 0.129
MonkeyOCR-pro-1.2B 87.0 0.084 85.0 84.2 89.0 0.130
Nanonets-OCR-s [17] 85.6 0.093 85.9 80.1 85.6 0.108
MinerU2-VLM [4] 85.6 0.078 81.0 83.5 87.7 0.086
olmOCR [23] 81.8 0.096 86.0 68.9 74.8 0.121
POINTS-Reader [14] 81.0 0.134 79.2 77.1 81.7 0.145
Mistral OCR [16] 78.8 0.164 82.8 70.0 78.0 0.144
OCRFlux [2] 74.8 0.193 68.0 75.8 80.2 0.202
Dolphin [7] 74.7 0.125 67.9 68.7 77.8 0.124
@@
Qwen3-VL-235B [1] 89.2 0.069 88.1 86.2 90.6 0.068
Gemini-2.5 Pro [8] 88.0 0.075 85.8 85.7 90.3 0.097
Qwen2.5-VL-72B [1] 87.0 0.094 88.3 82.2 86.2 0.102
InternVL3.5-241B [3] 82.7 0.142 87.2 75.0 81.3 0.125
InternVL3-78B 80.3 0.131 83.4 70.6 77.7 0.113
GPT-4o [19] 75.0 0.217 79.7 67.1 76.1 0.148
@@
PP-StructureV3 [6] 86.7 0.073 85.8 81.7 89.5 0.073
Mineru2-pipeline 75.5 0.209 76.6 70.9 79.1 0.225
Marker-1.8.2 [22] 71.3 0.206 76.7 57.9 71.2 0.250
Ours MonkeyOCR v1.5 92.9 0.046 91.2 92.0 95.0 0.049
@@ 3 Experiments
To evaluate our model’s capabilities in general document parsing and complex table recognition, we assess MonkeyOCR v1.5 on OmniDocBench [20] for overall performance, and further evaluate its table recognition on PubTabNet [32] and OCRFlux-pubtabnet-single [2]. As shown in Fig. 1, MonkeyOCR v1.5 demonstrates state-of-the-art performance in general document parsing and significantly outperforms the previous best model, PPOCR-VL, by 9.2% on complex table recognition tasks.
@@ 3.1 Comparison with Other Methods on Different Tasks
Document parsing involves multiple subtasks, including text recognition, formula recognition, table recognition, and reading-order prediction. Our model achieves the overall best performance, surpassing the previous state-of-the-art expert document parsing model, PPOCR-VL, by 1% on the overall metric, with improvements of 2.5% in formula recognition and 1% in table recognition. It also outperforms MinerU2.5 by 2.2% overall, including 2.7% in formula recognition and 3.8% in table recognition. Compared with pipeline-based approaches, our model exceeds the best pipeline method, PP-Struct v3, by 6.2%, demonstrating the effectiveness of our two-stage VLM-based paradigm. Against large general end-to-end models, our method outperforms Qwen-3-VL-235B by 3.7% and the best closed-source model, Gemini 2.5-Pro, by 4.9%, highlighting the advantage of specialized models in vertical domains.
@@
Marker-1.8.2  [22] 0.180 0.041 0.101 0.291 0.296 0.111 0.272 0.466 0.034
MinerU2-pipeline  [4] 0.424 0.023 0.263 0.122 0.082 0.395 0.074 0.260 0.041
$P$-StructureV3  [6] 0.079 0.024 0.042 0.111 0.095 0.072 0.062 0.124 0.018
@@
Qwen2.5-VL-72B [1] 0.042 0.080 0.059 0.115 0.068 0.096 0.238 0.123 0.026
Gemini-2.5 Pro  [8] 0.033 0.018 0.069 0.162 0.094 0.016 0.135 0.117 0.017
@@
Dolphin  [7] 0.096 0.045 0.062 0.133 0.168 0.070 0.239 0.256 0.019
OCRFlux  [2] 0.087 0.087 0.082 0.184 0.207 0.105 0.730 0.157 0.019
Mistral-OCR  [16] 0.092 0.053 0.061 0.135 0.134 0.058 0.564 0.310 0.052
POINTS-Reader  [14] 0.033 0.078 0.067 0.137 0.190 0.134 0.379 0.094 0.095
olmOCR-7B  [23] 0.050 0.037 0.054 0.120 0.073 0.070 0.292 0.122 0.046
@@
Nanonets-OCR-s  [17] 0.055 0.058 0.061 0.093 0.083 0.092 0.197 0.161 0.040
MonkeyOCR pro-1.2B  [13] 0.096 0.035 0.053 0.111 0.089 0.049 0.100 0.169 0.020
@@
dots.ocr  [25] 0.029 0.023 0.043 0.079 0.047 0.022 0.067 0.112 0.008
MinerU2.5 [18] 0.029 0.024 0.033 0.050 0.068 0.032 0.054 0.116 0.010
@@ 3.2 Comparison with Other Methods on Different Document Types
Dataset Nanonets-OCR[17] MonkeyOCR [13] OCRFlux[2] MinerU2.5[18] PaddleOCR-VL[5] Ours
@@ 3.3 Comparison with Other Methods on Table Recognition
To evaluate the performance of our model on table recognition tasks, we conduct experiments on PubTabNet and OCRFlux-pubtabnet-single, comparing MonkeyOCR v1.5 with several state-of-the-art methods, including MinerU2.5 and PaddleOCR-VL. As shown in Table 5, MonkeyOCR v1.5 achieves the best performance on both datasets. Notably, on the OCRFlux-complex subset, our model surpasses MinerU2.5 by 9.2%, demonstrating the effectiveness of our visual consistency-based reinforcement learning strategy in handling complex table layouts. These results confirm that MonkeyOCR v1.5 not only delivers superior recognition accuracy but also exhibits strong generalization to diverse and structurally intricate table formats commonly found in real-world documents.
Model Type Models Slides
Academic
Papers
Book Textbook Exam Papers Magazine Newspaper Notes
Financial
Report
Pipeline Tools Marker-1.8.2  [22] 0.180 0.041 0.101 0.291 0.296 0.111 0.272 0.466 0.034
MinerU2-pipeline  [4] 0.424 0.023 0.263 0.122 0.082 0.395 0.074 0.260 0.041
$P$-StructureV3  [6] 0.079 0.024 0.042 0.111 0.095 0.072 0.062 0.124 0.018
General VLMs GPT-4o  [19] 0.102 0.120 0.129 0.160 0.194 0.142 0.625 0.261 0.334
InternVL3-76B  [3] 0.035 0.105 0.063 0.083 0.101 0.041 0.583 0.092 0.067
InternVL3.5-24B 0.048 0.086 0.024 0.106 0.099 0.058 0.640 0.136 0.112
Qwen2.5-VL-72B [1] 0.042 0.080 0.059 0.115 0.068 0.096 0.238 0.123 0.026
Gemini-2.5 Pro  [8] 0.033 0.018 0.069 0.162 0.094 0.016 0.135 0.117 0.017
Specialized VLMs Dolphin  [7] 0.096 0.045 0.062 0.133 0.168 0.070 0.239 0.256 0.019
OCRFlux  [2] 0.087 0.087 0.082 0.184 0.207 0.105 0.730 0.157 0.019
Mistral-OCR  [16] 0.092 0.053 0.061 0.135 0.134 0.058 0.564 0.310 0.052
POINTS-Reader  [14] 0.033 0.078 0.067 0.137 0.190 0.134 0.379 0.094 0.095
olmOCR-7B  [23] 0.050 0.037 0.054 0.120 0.073 0.070 0.292 0.122 0.046
MinerU2-VLM 0.075 0.010 0.036 0.128 0.070 0.065 0.183 0.080 0.024
Nanonets-OCR-s  [17] 0.055 0.058 0.061 0.093 0.083 0.092 0.197 0.161 0.040
MonkeyOCR pro-1.2B  [13] 0.096 0.035 0.053 0.111 0.089 0.049 0.100 0.169 0.020
dotsOCR  [25] 0.029 0.023 0.043 0.079 0.047 0.022 0.067 0.112 0.008
MinerU2.5 [18] 0.029 0.024 0.033 0.050 0.068 0.032 0.054 0.116 0.010

### v2

@@ Abstract
Document parsing is a core task in document intelligence, supporting applications such as information extraction, retrieval-augmented generation, and automated document analysis. However, real-world documents often feature complex layouts with multi-level tables, embedded images or formulas, and cross-page structures, which remain challenging for existing OCR systems. We introduce MonkeyOCR v1.5, a unified vision–language framework that enhances both layout understanding and content recognition through a two-stage pipeline. The first stage employs a large multimodal model to jointly predict layout and reading order, leveraging visual information to ensure sequential consistency. The second stage performs localized recognition of text, formulas, and tables within detected regions, maintaining high visual fidelity while reducing error propagation. To address complex table structures, we propose a visual consistency–based reinforcement learning scheme that evaluates recognition quality via render-and-compare alignment, improving structural accuracy without manual annotations. Additionally, two specialized modules, Image-Decoupled Table Parsing and Type-Guided Table Merging, are introduced to enable reliable parsing of tables containing embedded images and reconstruction of tables crossing pages or columns. Comprehensive experiments on OmniDocBench v1.5 demonstrate that MonkeyOCR v1.5 achieves state-of-the-art performance, outperforming PPOCR-VL and MinerU 2.5 while showing exceptional robustness in visually complex document scenarios. A trial link can be found at github link https://github.com/Yuliang-Liu/MonkeyOCR.
@@ 1 Introduction
Traditional pipeline-based approaches [mineru; Cui2025PaddleOCR3T] break down document parsing into a series of subtasks, such as layout detection, text and formula detection, text recognition, and table or formula recognition, with each handled by a dedicated model. This multi-stage process is prone to error accumulation, where mistakes in earlier steps propagate and compromise overall performance. In contrast, end-to-end models [qwen2.5-vl; internvl; gpt4o; gemini25] process the entire document image in a single pass; however, the high resolution of document images produces a massive number of visual tokens, and the quadratic complexity of self-attention mechanisms creates a significant computational bottleneck. To address these limitations, MonkeyOCR [li2025monkeyocr] proposed the SRR paradigm, which decouples document parsing into structure detection, content recognition, and reading order prediction. This design streamlines the conventional multi-stage pipeline, effectively mitigating cumulative errors while avoiding the substantial computational overhead associated with full-page end-to-end processing, thereby advancing intelligent multimodal document understanding (Fig. 2). Mineru 2.5 [mineru2.5] further simplify the three-stage framework by employing a unified large multimodal model to jointly predict document layout and reading order, followed by content recognition. PPOCR-VL [paddleocrvl] adopts a similar three-stage methodology, utilizing lightweight models for structural analysis and reading order prediction, and subsequently applying a large multimodal model for content recognition.
@@
Experimental results show that MonkeyOCR v1.5 achieves state-of-the-art performance on the widely adopted OmniDocBench v1.5 benchmark, surpassing the previous best-performing methods PPOCR-VL and MinerU 2.5 by 0.15% and 2.34%, respectively, in overall performance. Notably, MonkeyOCR v1.5 demonstrates stronger robustness in complex scenarios. On newspaper documents characterized by dense text and intricate layouts, it achieves the best performance among all competitors. Moreover, on the OCRFlux-complex dataset, our method significantly outperforms the previous state-of-the-art PPOCR-VL by 8.2%. Beyond accuracy improvements, MonkeyOCR v1.5 further extends its capabilities over MonkeyOCR, enabling embedded image recovery, cross-page table reconstruction, and multi-column table merging, demonstrating strong potential in complex real-world document scenarios [ocrbench; ocrbenchv2].
@@
We propose MonkeyOCR v1.5, a vision-language-based document parsing framework designed for robust and efficient OCR in complex, real-world documents. Compared with the previous version, v1.5 simplifies the pipeline into two stages and introduces a visual-consistency-based reinforcement learning paradigm that improves table recognition accuracy without relying on large-scale manual annotations. In addition, to address challenges such as embedded images and cross-page or cross-column table merging that other methods struggle with (Tab. 1), we incorporate an image-decoupled and type-guided table recognition module. Together, these advancements make MonkeyOCR v1.5 a scalable and high-fidelity OCR system well-suited for heterogeneous document understanding.
@@
MinerU2.5 [mineru2.5] × × × × ×
MonkeyOCR [li2025monkeyocr] × × × × × × ×
OCRFlux [OCRFLUX] × × × × ×
dotsOCR [dotsocr] × × × × × × ×
PaddleOCR-VL [paddleocrvl] × × × × × ×
@@ 2.1 Overall Pipeline
The overall pipeline of MonkeyOCR v1.5 consists of two sequential yet lightweight stages: layout analysis and content recognition, as described in Fig. 3. This design leverages the model’s semantic understanding capabilities, enabling more accurate layout detection and reading order prediction.
@@ 2.3 Image-Decoupled Table Parsing
As illustrated in Fig. 5, we first employ YOLOv10 [yolov10] to detect embedded images within table regions. Each detected figure is replaced by a precisely sized placeholder mask, while a one-to-one mapping between placeholder IDs and cropped image files is maintained. The masked table is then passed to the recognizer, which is trained with an auxiliary objective encouraging it to treat placeholders as atomic tokens and to produce an intermediate HTML representation containing <img> tags at the corresponding locations. During post-processing, each <img> tag is deterministically replaced with its associated image according to the stored mapping, yielding a visually complete table while preserving clean textual outputs. This decoupled design effectively separates image restoration from structural recognition, improving both text accuracy and visual fidelity in complex tables.
@@ 2.4 Type-Guided Table Merging
PaddleOCR-VL  [paddleocrvl] 92.86 0.035 91.22 90.89 94.76 0.043
MinerU2.5  [mineru2.5] 90.67 0.047 88.46 88.22 92.38 0.044
MonkeyOCR-pro-3B  [li2025monkeyocr] 88.85 0.075 87.25 86.78 90.63 0.128
dots.ocr  [dotsocr] 88.41 0.048 83.22 86.78 90.62 0.053
Deepseek-OCR [deepseekocr] 87.01 0.073 83.37 84.97 88.80 0.086
MonkeyOCR-3B 87.13 0.075 87.45 81.39 85.92 0.129
MonkeyOCR-pro-1.2B 86.96 0.084 85.02 84.24 89.02 0.130
Nanonets-OCR-s  [Nanonets-OCR-s] 85.59 0.093 85.90 80.14 85.57 0.108
MinerU2-VLM  [2024mineru] 85.56 0.078 80.95 83.54 87.66 0.086
olmOCR  [olmocr] 81.79 0.096 86.04 68.92 74.77 0.121
POINTS-Reader  [points-reader] 80.98 0.134 79.20 77.13 81.66 0.145
Mistral OCR  [mistralocr] 78.83 0.164 82.84 70.03 78.04 0.144
OCRFlux  [OCRFLUX] 74.82 0.193 68.03 75.75 80.23 0.202
Dolphin  [Feng2025DolphinDI] 74.67 0.125 67.85 68.70 77.77 0.124
@@
Qwen3-VL-235B  [qwen2.5-vl] 89.15 0.069 88.14 86.21 90.55 0.068
Gemini-2.5 Pro  [gemini25] 88.03 0.075 85.82 85.71 90.29 0.097
Qwen2.5-VL-72B  [qwen2.5-vl] 87.02 0.094 88.27 82.15 86.22 0.102
InternVL3.5-241B  [internvl] 82.67 0.142 87.23 75.00 81.28 0.125
InternVL3-78B 80.33 0.131 83.42 70.64 77.74 0.113
GPT-4o  [gpt4o] 75.02 0.217 79.70 67.07 76.09 0.148
@@
$P$-StructureV3  [Cui2025PaddleOCR3T] 86.73 0.073 85.79 81.68 89.48 0.073
Mineru2-pipeline 75.51 0.209 76.55 70.90 79.11 0.225
Marker-1.8.2  [marker] 71.30 0.206 76.66 57.88 71.17 0.250
Ours MonkeyOCR v1.5 93.01 0.045 91.54 91.99 95.04 0.049
@@ 3 Experiments
To evaluate our model’s capabilities in general document parsing and complex table recognition, we assess MonkeyOCR v1.5 on OmniDocBench [omnidocbench] for overall performance, and further evaluate its table recognition on PubTabNet [pubtabnet] and OCRFlux-pubtabnet-single [OCRFLUX]. As shown in Fig. 1, MonkeyOCR v1.5 demonstrates state-of-the-art performance in general document parsing and significantly outperforms the previous best model, PPOCR-VL, by 9.2% on complex table recognition tasks.
@@ 3.1 Comparison with Other Methods on Different Tasks
Document parsing involves multiple subtasks, including text recognition, formula recognition, table recognition, and reading-order prediction. Our model achieves the overall best performance, surpassing the previous state-of-the-art expert document parsing model, PPOCR-VL, by 0.15% on the overall metric, with improvements of 0.32% in formula recognition and 1.01% in table recognition. It also outperforms MinerU2.5 by 2.34% overall, including 3.08% in formula recognition and 3.77% in table recognition. Compared with pipeline-based approaches, our model exceeds the best pipeline method, PP-Structure v3, by 6.38%, demonstrating the effectiveness of our two-stage VLM-based paradigm. Against large general end-to-end models, our method outperforms Qwen-3-VL-235B by 3.86% and the best closed-source model, Gemini 2.5-Pro, by 4.98%, highlighting the advantage of specialized models in vertical domains.
@@
Marker-1.8.2   [marker] 0.180 0.041 0.101 0.291 0.296 0.111 0.272 0.466 0.034
MinerU2-pipeline   [2024mineru] 0.424 0.023 0.263 0.122 0.082 0.395 0.074 0.260 0.041
$P$-StructureV3   [Cui2025PaddleOCR3T] 0.079 0.024 0.042 0.111 0.095 0.072 0.062 0.124 0.018
@@
Qwen2.5-VL-72B  [qwen2.5-vl] 0.042 0.080 0.059 0.115 0.068 0.096 0.238 0.123 0.026
Gemini-2.5 Pro   [gemini25] 0.033 0.018 0.069 0.162 0.094 0.016 0.135 0.117 0.017
@@
Dolphin   [Feng2025DolphinDI] 0.096 0.045 0.062 0.133 0.168 0.070 0.239 0.256 0.019
OCRFlux   [OCRFLUX] 0.087 0.087 0.082 0.184 0.207 0.105 0.730 0.157 0.019
Mistral-OCR   [mistralocr] 0.092 0.053 0.061 0.135 0.134 0.058 0.564 0.310 0.052
POINTS-Reader   [points-reader] 0.033 0.078 0.067 0.137 0.190 0.134 0.379 0.094 0.095
olmOCR-7B   [olmocr] 0.050 0.037 0.054 0.120 0.073 0.070 0.292 0.122 0.046
@@
Nanonets-OCR-s   [Nanonets-OCR-s] 0.055 0.058 0.061 0.093 0.083 0.092 0.197 0.161 0.040
MonkeyOCR pro-1.2B   [li2025monkeyocr] 0.096 0.035 0.053 0.111 0.089 0.049 0.100 0.169 0.020
@@
dots.ocr   [dotsocr] 0.029 0.023 0.043 0.079 0.047 0.022 0.067 0.112 0.008
MinerU2.5  [mineru2.5] 0.029 0.024 0.033 0.050 0.068 0.032 0.054 0.116 0.010
@@ 3.2 Comparison with Other Methods on Different Document Types
Dataset Nanonets-OCR [Nanonets-OCR-s] MonkeyOCR  [li2025monkeyocr] OCRFlux [OCRFLUX] MinerU2.5 [mineru2.5] PaddleOCR-VL [paddleocrvl] Ours
@@ 3.3 Comparison with Other Methods on Table Recognition
To evaluate the performance of our model on table recognition tasks, we conduct experiments on PubTabNet and OCRFlux-pubtabnet-single, comparing MonkeyOCR v1.5 with several state-of-the-art methods, including MinerU2.5 and PaddleOCR-VL. As shown in Table 5, MonkeyOCR v1.5 achieves the best performance on both datasets. Notably, on the OCRFlux-complex subset, our model surpasses PaddleOCR-VL by 9.2%, demonstrating the effectiveness of our visual consistency-based reinforcement learning strategy in handling complex table layouts. These results confirm that MonkeyOCR v1.5 not only delivers superior recognition accuracy but also exhibits strong generalization to diverse and structurally intricate table formats commonly found in real-world documents.